Body Exam measurements are from 300 subjects. Age is in years, GENDER 1 = Male 0 = Female. PULSE is the pulse rate (beats per minute), SYSTOTOLIC is systolic blood pressure (mmHg). DIASTOLIC is diastolic blood pressure (mmHG), HDL is cholesterol (mg/dL). LDL is LDL cholesterol (Hg/dL). WHITE is the white blood cell count (1000cells/μL). RED is red blood cell count (million cells/ μL), PLATE is platelet count (1000cells/μL). WEIGHT is weight (kg), HEIGHT is height (cm), WAIST is waist circumference (cm), ARM CIRC is arm circumference (cm), and BMI is body mass index (kg/m²).

GENDER (1=M)   PULSE
0   36
1   40
1   42
0   44
1   50
1   50
0   52
1   52
1   52
1   52
1   52
1   52
0   54
0   54
1   54
1   54
1   54
1   54
1   54
1   54
0   56
0   56
0   56
0   56
0   56
0   56
0   56
0   56
0   56
1   56
1   56
0   58
0   58
0   58
0   58
0   58
1   58
1   58
1   58
1   58
1   58
1   58
1   58
1   58
0   60
0   60
0   60
1   60
1   60
1   60
1   60
1   60
1   60
1   60
0   62
0   62
0   62
0   62
0   62
0   62
0   62
0   62
1   62
1   62
1   62
1   62
1   62
1   62
1   62
1   62
1   62
1   62
1   62
0   64
0   64
0   64
0   64
0   64
0   64
0   64
0   64
0   64
1   64
1   64
1   64
1   64
1   64
1   64
1   64
1   64
1   64
1   64
1   64
1   64
1   64
0   66
0   66
0   66
0   66
0   66
0   66
0   66
0   66
1   66
1   66
1   66
1   66
1   66
1   66
1   66
1   66
1   66
1   66
1   66
1   66
1   66
1   66
1   66
1   66
1   66
0   68
0   68
0   68
0   68
0   68
1   68
1   68
1   68
1   68
1   68
0   70
0   70
0   70
0   70
0   70
1   70
1   70
1   70
1   70
1   70
1   70
1   70
1   70
0   72
0   72
0   72
0   72
0   72
0   72
0   72
0   72
0   72
0   72
0   72
1   72
1   72
1   72
1   72
1   72
1   72
1   72
1   72
1   72
1   72
1   72
1   72
1   72
1   72
0   74
0   74
0   74
0   74
0   74
0   74
0   74
0   74
0   74
0   74
0   74
1   74
1   74
1   74
1   74
1   74
1   74
1   74
1   74
1   74
0   76
0   76
0   76
0   76
0   76
0   76
0   76
1   76
1   76
1   76
1   76
1   76
1   76
0   78
0   78
0   78
0   78
0   78
0   78
0   78
1   78
1   78
1   78
1   78
1   78
1   78
0   80
0   80
0   80
0   80
0   80
0   80
0   80
0   80
0   80
0   80
1   80
1   80
1   80
1   80
1   80
1   80
0   82
0   82
0   82
0   82
0   82
0   82
0   82
0   82
0   82
0   82
0   82
1   82
1   82
1   82
1   82
1   82
1   82
1   82
1   82
0   84
0   84
0   84
0   84
0   84
0   84
0   84
0   84
1   84
1   84
1   84
1   84
1   84
0   86
0   86
0   86
0   86
0   86
1   86
1   86
1   86
1   86
0   88
0   88
0   88
1   88
1   88
0   90
0   90
0   90
0   90
1   90
1   92
0   94
0   94
0   94
0   94
1   94
0   96
0   96
0   96
1   96
1   96
0   98
0   98
0   100
0   100
1   102
0   104
0   104
1   104

What is the probability that a randomly selected individual has a pulse rate of 58?

What is the probability that a randomly selected individual has a pulse rate of 58? Given that the individual is a female?

Two individuals are randomly selected from the data without replacement, find the probability that the first person selected is a female and the second person selected is a male?
Answer #1

Total entries = 300

no. of entries with 58 pulse rate = 13

so, probability that a randomly selected individual has a pulse rate of 58 = 13/300 = 0.0433

no. of females having 58 pulse rate = 5

so, probability that a randomly selected individual given a female has a pulse rate of 58 = 5/300 = 0.0167

no. of females = 144

no. of males = 156

probability first person is female = 144/300

probability second person is male = 156/299

so total probability = =0.25